OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1) @DanJS

OpenStudy (jackthegreatest):

for x and y?

OpenStudy (alexandervonhumboldt2):

x=14+3y x=9-3y

OpenStudy (danjs):

yeah, tons of people responding,, ill sit this one out , less you need help after

OpenStudy (alexandervonhumboldt2):

14+2y=9-3y

OpenStudy (alexandervonhumboldt2):

14-9=y

OpenStudy (anonymous):

I assume so @jackthegreatest

OpenStudy (anonymous):

plug in the numbers and solve the equation, PEMDAS. the first number for x and the second number for y

OpenStudy (jackthegreatest):

k

OpenStudy (alexandervonhumboldt2):

i mistaked somewhere

OpenStudy (alexandervonhumboldt2):

14+3y=9-3y

OpenStudy (anonymous):

Yeah I was kinda confused because you had (4,5) but those aren't options.

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

oh no

OpenStudy (alexandervonhumboldt2):

i got confused

OpenStudy (anonymous):

That's okay we all do @AlexandervonHumboldt2

OpenStudy (danjs):

slow down there alex, it is not a race for medals, well maybe it is for you

OpenStudy (anonymous):

I don't know where to start..

OpenStudy (anonymous):

First of all, this is my first time ever being on here. This is only my second question so I don't even know how to give medals. I mean unless someone told me how I wouldn't be able to...

OpenStudy (anonymous):

Can someone just help me figure it out..

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (danjs):

x - 2y = 14 x + 3y = 9

OpenStudy (danjs):

Multiply the first equation through by -1. (-1)(x - 2y) = (-1) *14 -x + 2y = -14 all the signes switched

OpenStudy (danjs):

now the two equations are -x + 2y = -14 x + 3y = 9

OpenStudy (danjs):

Add them together, Notice the X cancels out -x + 2y = -14 x + 3y = 9 --------------------- 0x + 5y = -5

OpenStudy (danjs):

You are left with 5y = -5 divide by 5 on both sides y = -1

OpenStudy (danjs):

you really don't have to figure out X in this problem, only one answer choice has y=-1 in (x , y) ---- (x , -1)

OpenStudy (anonymous):

Well two of the options has -1.. there's B and C

OpenStudy (danjs):

Each of those answers is a point in the form (x , y) the first spot is the x value, the second spot is the Y value. We found the y value, y=-1, only one answer has -1 in the Y spot

OpenStudy (anonymous):

unless it's just C

OpenStudy (danjs):

right , the third one (12 , -1) ...x=12, y = -1

OpenStudy (anonymous):

Okay okay that makes sense